OpenStudy (dls):

A smaller sphere is removed from a larger sphere.The sphere is an insulator of charge "q". We have to find the the electric field due to remaining portion at point A.

OpenStudy (dls):



OpenStudy (dls):

@yrelhan4 @electrokid

OpenStudy (anonymous):

this is a repeated problem..

OpenStudy (anonymous):

find \(\sigma\)=charge per unit area when the smaller circel is removed, the charge on the Area gets removed then the same old

OpenStudy (dls):

Can we solve this with negative mass concept?

OpenStudy (dls):

 I kept back the removed thing..i want to solve this way

OpenStudy (anonymous):

let know if you got lost in it

OpenStudy (dls):

That is for electric field in the cavity,i want remaining thing

OpenStudy (dls):

@JFraser

OpenStudy (anonymous):

for outside, it does not matter the sphere now acts like a point charge but with a smaller total charge!!

OpenStudy (anonymous):

\[ \sigma={q\over4\pi 4R^2}={q\over16\pi R^2}\\ Q_{\rm left}=\sigma \times\left(16\pi R^2-4\pi R^2\right)={q\over16\pi R^2}\times12\pi R^2\\ Q_{\rm left}={3q\over 4}\\ E_A=K{Q_{\rm left}\over d_A^2} \]

OpenStudy (vincent-lyon.fr):

Answer will show that electric field is uniform in the cavity. You can solve by superposition of positive and negative charges, no problem, since laws of electromagnetism are all linear.

OpenStudy (anonymous):

@Vincent-Lyon.Fr not In the cavity.. outside the system at a point "A"